OpenStudy (anonymous):

What is the slope-intercept form of the equation y – 7 = –5/2(x + 4)? (1 point)

OpenStudy (hoblos):

y – 7 = –5/2(x + 4) y= (-5/2)x -10 +7 y=(-5/2)x -3

OpenStudy (anonymous):

http://mathway.com/problem.aspx?p=precalculus try this site:"DD